# On the Compatibility of Constructive Predicative Mathematics with Weyl’s Classical Predicativity

### v2

@@ Abstract
It is well known that most constructive and predicative foundations aiming to develop Bishop’s constructive analysis are incompatible with a classical predicative development of analysis as Weyl put forward in his Das Kontinuum.
Here, we show how this incompatibility arises from the possibility to define sets by quantifying over (the exponentiation of) functional relations. Such a possibility is present in most constructive foundations but it is not allowed in modern reformulations of Weyl’s logical system. In particular, we argue how in Aczel’s Constructive Set Theory, Martin-Löf’s type theory and Homotopy Type Theory, the incompatibility with classical predicativity à la Weyl reduces to the fact of being able to interpret Heyting arithmetic in all finite types with the addition of the internal rule of number-theoretic unique choice, identifying functional relations over natural numbers with a primitive notion of function defined as $\lambda$-terms of type theory.
Then, we argue how a possible way out is offered by constructive foundations, such as the Minimalist Foundation, where exponentiation is limited to functions defined as $\lambda$-terms of (dependent) type theory.
@@ 1 Introduction
The price to be paid in MF  is to renounce the so-called rule of unique choice identifying functional relations with $\lambda$-terms. In this way, exponentiation of type-theoretic functions does not entail exponentiation of functional relations as it happens instead in Martin-Löf’s type theory and in Homotopy Type Theory, whose logics are also sorted on a dependent type theory. Accordingly, dropping the rule of unique choice entails dropping all the number-theoretic choice principles characteristic of Bishop’s mathematics as well as the fact that real numbers à la Dedekind or Cauchy form a set.
@@
The structure of the paper is the following: first, we give a short overview of Weyl’s classical predicative conception in Das Kontinuum and highlight the points of convergence with some ideas underlying the Minimalist Foundation. In Section 3 we present some well-known examples of constructive predicative foundations. Then, in Section 4 we show that already Heyting Arithmetic in all finite types is incompatible with classical predicativity if it is extended with classical logic and the (number-theoretic) internal rule of unique choice. In Section 5 we extend this argument to other systems and analyze the incompatibility of MLTT, HoTT and CZF with classical predicativity. Finally, in Section 6 we introduce the Minimalist Foundation and we argue that it might represent a natural solution to the issues discussed in the previous sections.
@@ 2 Weyl’s Classical Predicativity
There have been various attempts in the literature to formalize Weyl’s system in a modern fashion since Weyl’s original presentation presents some ambiguities. An influential reconstruction of Weyl’s system has been carried on by Feferman in several works [16, 17, 18]. Feferman has extensively argued that Weyl’s system should correspond to the theory $\mathbf{ACA}_{0}$ or, if the full induction scheme is accepted, to $\mathbf{ACA}$, which are well-known subsystems of second-order arithmetic [48]. But such a reconstruction has been criticized by Adams and Luo in [4, 3], and more recently by Avron in [5]. They observed that in certain sections of Das Kontinuum, Weyl introduces non-arithmetical sets and makes use of induction for non-arithmetical formulas. Therefore, they concluded that Weyl’s system should be stronger than Feferman’s reformulation of Weyl’s system, even if ${\bf ACA}_{0}$ should be sufficient to reconstruct predicative analysis as developed in Das Kontinuum.
The systems introduced in these papers are many-sorted logics axiomatized over a dependent type theory, in a way similar to both levels of the Minimalist Foundation. They all differ in the way exponentiation of functions is treated. Indeed, Adams and Luo’s system in [4, 3] includes some exponentiation of $\lambda$-functions in the form of function types, which are treated as large types in the sense that they are not contained in the universe of small types, that in turn can be thought of as the universe of sets. Instead, Avron in [5] restricts the possibility of forming exponentials by assuming that only certain types might serve as codomains by not allowing to define sets by quantification over functions. On the contrary, in MF there is no exponentiation of functions between collections but only sets exponentiating $\lambda$-functions between sets.
@@ 3 Constructive Predicative Theories
Among the most popular frameworks introduced as foundations for Bishop’s constructive mathematics there are Constructive Zermelo-Fraenkel set theory $\mathbf{CZF}$ [1] and Martin-Löf Type Theory MLTT  [40]. More recently, an extension of the latter called Homotopy Type Theory HoTT [53] has gained much attention, especially for its applications to synthetic homotopy theory and higher categories. All these systems are examples of constructive and predicative theories. However, the predicativity of these theories is closely tied to the adoption of a constructive logic. Indeed, whenever they are extended with the law of excluded middle or equivalent classical principles, they become impredicative and thus they are not compatible with classical predicativity a lá Weyl. Briefly, the issue is that such theories are closed under the exponentiation of functional relations and the two-element set that, when classical logic is added, are strong enough to allow for impredicative constructions. Therefore, we cannot establish the predicativity of their classical counterparts, as it happens instead in the case of Heyting arithmetic or the Minimalist Foundation, which we will introduce in detail in Section 6.
@@ 3.2 Martin-Löf’s type theory
Types of MLTT are introduced through the following kinds of rules: formation, introduction, elimination, and computation. In particular, the constructors of a given type are specified by the introduction rules, while the elimination rule states simultaneously a principle of induction at the level of types and of definition by recursion at the level of terms with the aid of the computation rules. The predicative nature of the theory is related to the fact that all the types, also called sets, are inductively generated by the relevant rules. Two fundamental aspects of MLTT are the presence of a primitive notion of function, called here $\lambda$-function to disambiguate it from the notion of function as functional relation, and the identification between propositions and types. $\lambda$-functions are represented as $\lambda$-terms and this differentiates this foundational theory from other theories such as CZF. The identification between propositions and types makes the constructive logic of MLTT stronger than usual intuitionistic logic. This additional strength is visible in the case of the existential quantifier that is identified with the dependent sum type $\Sigma$. The elimination rule for this type is enough strong to allow for the definition of projection operators [^2] We recall that the canonical elements of $\Sigma$-types given by the introduction rule are pairs.. These projections play a fundamental role in establishing the following fact:
@@
We remark that the rules for $\Sigma$-type are enough strong to derive also the corresponding choice rule. Furthermore, the Axiom of Unique Choice AC!, which identifies functional relations with $\lambda$-functions,
@@ 6 A possible way out: the Minimalist Foundation
We call number-theoretic internal rule of unique choice the instance of the internal rule of unique choice when $A$ and $B$ coincide with the set $\mathbb{N}$ of natural numbers.
@@
$({\bf AC!_{\mathbb{N},\mathbb{N}}})\ \forall_{x\in\mathsf{N}}\ \exists!_{y\in\mathbb{N}}\ R(x,y)\ \rightarrow\ \exists_{f\in\mathbb{N}\rightarrow\mathbb{N}}\ \forall_{x\in\mathbb{N}}\ R(x,f(x))$
@@
First, observe that $\bf AC!_{\mathbb{N},\mathbb{N}}$ implies the number-theoretic internal rule of unique choice. The claim then follows from Proposition 6.7 and Cor. 4.4. ∎
Since the validity of the axiom of unique choice in a theory can be transferred to its extensions, contrary to the rule of unique choice, from Cor. 6.8 we can immediately deduce the non-validity of ${\bf AC!_{\mathbb{N},\mathbb{N}}}$ in mTT and emTT themselves:
@@
It is sufficient to define a translation from ACA to emTT as follows. First-order variables of ACA are translated as variables of type $\mathbb{N}$ in emTT. The constants for zero and the successor are interpreted by the corresponding constructors of type $\mathbb{N}$ in emTT. Second-order variables for sets of natural numbers are interpreted as propositional functions $X\in\mathbb{N}\rightarrow\mathcal{P}(1)$ in emTT and the atomic formulas of the form $t\in X$ are translated as applications of these functions $X(t)$.
In general, first-order formulas of ACA are interpreted as small propositions of emTT with quantifiers and equalities restricted to $\mathbb{N}$. While second-order formulas are interpreted by propositions of emTT quantifying over collections.
Peano axioms can be derived from the rules for natural numbers in emTT and the presence of power-collections, see [24]. The induction scheme of ACA follows from the elimination rule for $\mathsf{N}$ which acts towards all types including all propositions. Finally, the arithmetical comprehension axiom is validated as follows: given an arithmetical formula $\phi(x)$ in ACA, we know that it is interpreted as a small proposition in emTT. Then we can introduce $[\phi](x)\in[x\in\mathbb{N}]$ in $\mathcal{P}(1)$ and we get $X\ {\ :\equiv\ }\lambda x.\phi(x)\in\mathbb{N}\rightarrow\mathcal{P}(1)$. Therefore, it follows that there exists $X$ such that for all $n\in\mathbb{N}$, $X(x)\leftrightarrow\phi$. ∎
@@
A direct proof of this was given [26] but it follows also from the predicativity of the classical version of MF (see also [33]). If real numbers form a set in MF then this is so in the classical version of MF and every interval of real numbers form a set too. Since classically, the power-collection of natural numbers can be put in bijection with an interval of real numbers, then it would follow that impredicative definitions can be employed in the classical version of MF contrary to its predicativity established thanks in prop.6.7. ∎
@@
Therefore, one can carry on in MF the development of analysis initiated by Palmgren in [7, 42, 43] (see also [21]) and that of Positive Topology in [47] and loc.cit..
@@ 7 Conclusion
In the future, besides establishing the exact proof-theoretic strength of MF, which is currently an open problem, we will explore whether the extensions of MF  with inductive and coinductive definitions in [28, 29, 32] are still equiconsistent with their classical counterpart, as established for MF in [33], or at least they are still predicative after the addition of the law of excluded middle. A positive answer to either of these questions would provide a more expressive constructive foundation capable of formalizing general results of point-free mathematics as those presented in [12, 37, 15, 47], by keeping the compatibility both with Bishop’s constructivism and with Weyl’s classical predicativity, as in the case of MF.

### v3

@@ Abstract
It is well known that most constructive and predicative foundations aiming to develop Bishop’s constructive analysis are incompatible with a classical predicative development of analysis as put forward by Weyl in his Das Kontinuum.
Here, we show how this incompatibility arises from the possibility to define sets by quantifying over (the exponentiation of) functional relations. Such a possibility is present in most constructive foundations but it is not allowed in modern reformulations of Weyl’s logical system. In particular, we show how in Aczel’s Constructive Set Theory, Martin-Löf’s type theory and Homotopy Type Theory, the incompatibility with classical predicativity à la Weyl reduces to the fact of being able to interpret Heyting arithmetic in all finite types with the addition of the internal rule of number-theoretic unique choice, identifying functional relations over natural numbers with a primitive notion of function defined as $\lambda$-terms of type theory.
Then, we argue that a possible way out is offered by constructive foundations, such as the Minimalist Foundation, where exponentiation is limited to functions defined as $\lambda$-terms of (dependent) type theory.
@@ 1 Introduction
The price to be paid in MF  is giving up the so-called rule of unique choice identifying functional relations with $\lambda$-terms. In this way, exponentiation of type-theoretic functions does not entail exponentiation of functional relations as it happens instead in Martin-Löf’s type theory and in Homotopy Type Theory, whose logics are also sorted on a dependent type theory. Accordingly, dropping the rule of unique choice entails dropping all the number-theoretic choice principles characteristic of Bishop’s mathematics as well as the fact that real numbers à la Dedekind or Cauchy form a set.
@@
The structure of the paper is the following: first, we give a short overview of Weyl’s classical predicative conception in Das Kontinuum and highlight the points of convergence with some ideas underlying the Minimalist Foundation. In Section 3 we present some well-known examples of constructive predicative foundations. Then, in Section 4 we show that already Heyting Arithmetic in all finite types is incompatible with classical predicativity if it is extended with classical logic and the (number-theoretic) internal rule of unique choice. In Section 5 we extend this argument to other systems and analyze the incompatibility of MLTT, HoTT and CZF with classical predicativity. Finally, in Section 6 we introduce the Minimalist Foundation and argue that it might represent a natural solution to the issues discussed in the previous sections.
@@ 2 Weyl’s Classical Predicativity
There have been various attempts in the literature to formalize Weyl’s system in a modern fashion, since Weyl’s original presentation presents some ambiguities. An influential reconstruction of Weyl’s system has been carried out by Feferman in several works [16, 17, 18]. Feferman has extensively argued that Weyl’s system should correspond to the theory $\mathbf{ACA}_{0}$ or, if the full induction scheme is accepted, to $\mathbf{ACA}$, which are well-known subsystems of second-order arithmetic [48]. But such a reconstruction has been criticized by Adams and Luo in [4, 3], and more recently by Avron in [5]. They observed that in certain sections of Das Kontinuum, Weyl introduces non-arithmetical sets and makes use of induction for non-arithmetical formulas. Therefore, they concluded that Weyl’s system should be stronger than Feferman’s reformulation of Weyl’s system, even if ${\bf ACA}_{0}$ should be sufficient to reconstruct predicative analysis as developed in Das Kontinuum.
The systems introduced in the above-mentioned papers are many-sorted logics axiomatized over a dependent type theory, in a way similar to both levels of the Minimalist Foundation. They all differ in the way exponentiation of functions is treated. Indeed, Adams and Luo’s system in [4, 3] includes some exponentiation of $\lambda$-functions in the form of function types, which are treated as large types in the sense that they are not contained in the universe of small types, that in turn can be thought of as the universe of sets. Instead, Avron in [5] restricts the possibility of forming exponentials by assuming that only certain types might serve as codomains by not allowing to define sets by quantification over functions. On the contrary, in MF there is no exponentiation of functions between collections but only sets exponentiating $\lambda$-functions between sets.
@@ 3 Constructive Predicative Theories
Among the most popular frameworks introduced as foundations for Bishop’s constructive mathematics there are Constructive Zermelo-Fraenkel set theory $\mathbf{CZF}$ [1] and Martin-Löf Type Theory MLTT  [40]. More recently, an extension of the latter called Homotopy Type Theory HoTT [53] has gained much attention, especially for its applications to synthetic homotopy theory and higher categories. All these systems are examples of constructive and predicative theories. However, the predicativity of these theories is closely tied to the adoption of a constructive logic. Indeed, whenever they are extended with the law of excluded middle or equivalent classical principles, they become impredicative and thus they are not compatible with classical predicativity à la Weyl. Briefly, the issue is that such theories are closed under the exponentiation of functional relations and the two-element set that, when classical logic is added, are strong enough to allow for impredicative constructions. Therefore, we cannot establish the predicativity of their classical counterparts, as it happens instead in the case of Heyting arithmetic or the Minimalist Foundation, which we will introduce in detail in Section 6.
@@ 3.2 Martin-Löf’s type theory
Types of MLTT are introduced through the following kinds of rules: formation, introduction, elimination, and computation. In particular, the constructors of a given type are specified by the introduction rules, while the elimination rule states simultaneously a principle of induction at the level of types and of definition by recursion at the level of terms with the aid of the computation rules. The predicative nature of the theory is related to the fact that all the types, also called sets, are inductively generated by the relevant rules. Two fundamental aspects of MLTT are the presence of a primitive notion of function, called here $\lambda$-function to disambiguate it from the notion of function as functional relation, and the identification between propositions and types. $\lambda$-functions are represented as $\lambda$-terms and this differentiates this foundational theory from other theories such as CZF. The identification between propositions and types makes the constructive logic of MLTT stronger than usual intuitionistic logic. This additional strength is visible in the case of the existential quantifier that is identified with the dependent sum type $\Sigma$. The elimination rule for this type is strong enough to allow for the definition of projection operators [^2] We recall that the canonical elements of $\Sigma$-types given by the introduction rule are pairs.. These projections play a fundamental role in establishing the following fact:
@@
We remark that the rules for $\Sigma$-type are strong enough to derive also the corresponding choice rule. Furthermore, the Axiom of Unique Choice AC!, which identifies functional relations with $\lambda$-functions,
@@ 6 A possible way out: the Minimalist Foundation
We call number-theoretic internal rule of unique choice the instance of the internal rule of unique choice when $A$ and $B$ coincide with the set $\mathsf{N}$ of natural numbers.
@@
$({\bf AC!_{\mathsf{N},\mathsf{N}}})\ \forall_{x\in\mathsf{N}}\ \exists!_{y\in\mathsf{N}}\ R(x,y)\ \rightarrow\ \exists_{f\in\mathsf{N}\rightarrow\mathsf{N}}\ \forall_{x\in\mathsf{N}}\ R(x,f(x))$
@@
First, observe that $\bf AC!_{\mathsf{N},\mathsf{N}}$ implies the number-theoretic internal rule of unique choice. The claim then follows from Proposition 6.7 and Cor. 4.4. ∎
Since the validity of the axiom of unique choice in a theory can be transferred to its extensions, contrary to the rule of unique choice, from Cor. 6.8 we can immediately deduce the non-validity of ${\bf AC!_{\mathsf{N},\mathsf{N}}}$ in mTT and emTT themselves:
@@
It is sufficient to define a translation from ACA to emTT as follows. First-order variables of ACA are translated as variables of type $\mathsf{N}$ in emTT. The constants for zero and the successor are interpreted by the corresponding constructors of type $\mathsf{N}$ in emTT. Second-order variables for sets of natural numbers are interpreted as propositional functions $X\in\mathsf{N}\rightarrow\mathcal{P}(1)$ in emTT and the atomic formulas of the form $t\in X$ are translated as applications of these functions $X(t)$.
In general, first-order formulas of ACA are interpreted as small propositions of emTT with quantifiers and equalities restricted to $\mathsf{N}$. While second-order formulas are interpreted by propositions of emTT quantifying over collections.
Peano axioms can be derived from the rules for natural numbers in emTT and the presence of power-collections, see [24]. The induction scheme of ACA follows from the elimination rule for $\mathsf{N}$ which acts towards all types including all propositions. Finally, the arithmetical comprehension axiom is validated as follows: given an arithmetical formula $\phi(x)$ in ACA, we know that it is interpreted as a small proposition in emTT. Then we can introduce $[\phi](x)\in[x\in\mathsf{N}]$ in $\mathcal{P}(1)$ and we get $X\ {\ :\equiv\ }\lambda x.\phi(x)\in\mathsf{N}\rightarrow\mathcal{P}(1)$. Therefore, it follows that there exists $X$ such that for all $n\in\mathsf{N}$, $X(x)\leftrightarrow\phi$. ∎
@@
A direct proof of this was given in [26] but it follows also from the predicativity of the classical version of MF (see also [33]). If real numbers form a set in MF then this is so in the classical version of MF and every interval of real numbers form a set too. Since classically, the power-collection of natural numbers can be put in bijection with an interval of real numbers, then it would follow that impredicative definitions can be employed in the classical version of MF contrary to its predicativity established thanks in prop.6.7. ∎
@@
Therefore, one can carry on in MF the development of analysis as pursued in [7, 42, 43] (see also [21]) and that of Positive Topology in [47] and loc.cit..
@@ 7 Conclusion
In the future, besides establishing the exact proof-theoretic strength of MF, which is currently an open problem, we will explore whether the extensions of MF  with inductive and coinductive definitions in [28, 29, 32] are still equiconsistent with their classical counterpart, as established for MF in [33], or at least whether they remain predicative after the addition of the law of excluded middle. A positive answer to either of these questions would provide a more expressive constructive foundation capable of formalizing general results of point-free mathematics as those presented in [12, 37, 15, 47], by keeping the compatibility both with Bishop’s constructivism and with Weyl’s classical predicativity, as in the case of MF.